OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>?

OpenStudy (surry99):

use dot product.

OpenStudy (anonymous):

Okay, so 2*3=6 -4*-8= 32 and 6+32 =38?

OpenStudy (anonymous):

but thats not one of my answer choices @surry99

OpenStudy (anonymous):

My choices are: 3.0° 6.0° -7.0° 16.0°

OpenStudy (surry99):

dot product says: u.v = |u||v|cos(theta)

OpenStudy (anonymous):

oh, where do you get cos theta thought?

OpenStudy (anonymous):

though*

OpenStudy (surry99):

theta is the angle between the vectors by defintion

OpenStudy (anonymous):

hmmm, okay.. I don't know how to set that up. U and V are 6 and 32 though, right?

OpenStudy (anonymous):

Oh! I just figured it out, nevermind(:

OpenStudy (surry99):

you did the left hand side...u.v now you need the: magnitude of u = |u| magnitude of v = |v|

OpenStudy (surry99):

good

OpenStudy (anonymous):

thank you! @surry99

OpenStudy (surry99):

you are welcome!